OpenStudy (anonymous):

factor -x2+6x+16 in polynomial descending order

OpenStudy (anonymous):

is it -x(x+8)(x+2)

OpenStudy (cwrw238):

no - if you expand that it includes -x^3

OpenStudy (cwrw238):

-x2+6x+16 = -(x^2 - 6x - 16) = (x + 2)(x - 8)